[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Wed Jan 18, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Wed Jan 18 19:27:12 UTC 2017


TAFB Shift Log for Wed Jan 18 2017 14:27:05 GMT-0500 (EST)

Shift: Day Shift

Pacific Forecaster: C McElroy

Atlantic Forecaster: C McElroy

Surface Analyst: M Formosa

Technical Issues Summary:

GFE/NCP issues: 0700L - Chris Mello came through to make sure all AWIPS  
terminals were logged off for the AWIPS upgrade. CAM 1000L - AWIPS upgrade  
complete but diagnostics are underway to check out the build. CAM 1115L -  
Briefed Henry at HFO on progress of AWIPS upgrrade; told him it looks like  
HFO will not have to do any more products but am still waiting for go ahead  
to take back EPAC products. CAM 1139L - Chris Mello gave the go ahead to  
take back products. Called HFO to take back EPAC products. 1150L - Both  
EPAC and ATLC desks have reassumed product responsibility. 1325L - All  
graphical products from the 12Z cycle are on the web.

N-AWIPS issues:

Other issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170118/66c27d23/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list